OpenStudy (anonymous):

A wire is first bent into the shape of a triangle. Each side of the triangle is 16 inches long. Then the wire is unbent and reshaped into a square. What is the length of a side of the square?

OpenStudy (anonymous):

@iheartfood

OpenStudy (whpalmer4):

How long is the wire, if it makes a triangle with 3 equal sides of 16 inches?

OpenStudy (anonymous):

each side is 12 inches

OpenStudy (anonymous):

16•3=48 and 48/4=12

OpenStudy (anonymous):

ok its 12 inches because 16 times 3 is 48. 48 divided by 4 is 12

OpenStudy (anonymous):

Yes what @LadyKatsaTheGraceling said is correct.. you're looking for the perimeter of the triangle, so each side is 16, 3 sides = 16 * 3=48 and when it's unbent, there are still a total of 48 inches. 4 congruent (? lol i think that's the term:P ) sides in a square. so 48/4=12 each side is 12 inches long. voila!